A telemarketer makes six phone calls per hour and is able to make a sale on 30 percent of
these contacts. During the next two hours, find:
a. The probability of making exactly four sales.
b. The probability of making no sales.
c. The probability of making exactly two sales.
d. The mean number of sales in the two-hour period.

Answer by ikleyn(52781)

.
A telemarketer makes six phone calls per hour and is able to make a sale
on 30% of these contacts. During the next two hours. Find:
~~~~~~~~~~~~~~~~~~~~~~~~~~


Assuming that the next 2 hours he/she works at the same intensity
(does not have a coffee break/(lunch break) and continues his working hours),
in the next two hours, he/she makes 12 phone calls.

a) The probability of exactly making four sales 

   Ans: 12C4*0.3^4*0.7^(12-4) = 495*0.3^4*0.7^8 = 0.2311 (rounded).



b) The probability of making no sales.
   Ans: 0.7^12 = 0.01384 (rounded).



c) The probability of exactly making two sales.
   Ans: 12C2*0.3^2*0.7^(12-2) = 66*0.3^2*0.7^10 = 0.1678  (rounded).



d) The mean number of sales within the two-hour period.
   mean = np = 12*0.3 = 3.6 sales

Solved.